if 420(tanA+cotA)=841 then what is value of (116cosA - 58sinA)

Grade:

1 Answers

Rinkoo Gupta
10 years ago
420(tanA+cotA)=841
420(sinA/cosA +cosA/sinA) =841
420/sinAcosA=841
sin2A=840/841
cos2A=sqrt of (1681/841)=41/841
cosA=sqrt of (1+cos2A)/2=21/29
sinA=sqrt of (1-cos2A)/2=20/29

116cosA-58sinA=116(21/29)-58(20/29)=44
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty